INVESTIGATION INTO THE SENSITIVITY OF BEAM-BREAKUP INSTABILITY THRESHOLDS TO BEAMLINE COMPONENT MISALIGNMENTS

FOR THE CORNELL ENERGY RECOVERY LINAC

J. A. Crittenden, ^{∗} G.H. Hoffstaetter, M. Liepe, C.E. Mayes, D.C. Sagan, and N.R.A. Valles CLASSE, Cornell University, Ithaca, NY 14853-8001

Extensive lattice design studies have been performed for the proposed Cornell Energy Recovery Linac, including the level of orbit and optics distortions expected from a wide range of beamline component field errors and misalignments. In addition, beam-breakup calculation algorithms have been developed in the general framework of the Cornell X-ray ERL design software, enabling their extension to multi-pass optics design for ERLs. We report on investigations into the sensitivity of the calculated instability thresholds to a variety of beamline misalignments.

∗crittenden@cornell.edu